If the average power density for large scale wind turbine is 2 W/m2 , what will be the area required for the wind turbines with total energy of 4 TW ( 1 TW = 1012 W , that is, one terawatt is equal to one trillion watt;1kilometer=1000meter

Question 42 options:

1 million (Km)2

1.5 million (Km)2

1.7 million (Km)2

2.0 million (Km)2

2.2 million (Km)2
Answer #1

We see that area required for wind turbines will be given as equal to=4*10^12/2=2*10^12 square meter=2*10^6 square km=2 million (Km)^2